OpenStudy (arabpride):

Alrighty!!!!!~ Last QUESTION on my geom. review quiz Can I get some help? *^-^ 40minutes left though cx

OpenStudy (arabpride):

The two cones are similar. Find the volume of the larger cone. Round your answer to the nearest hundredth. A.) 2,025.73 ft^3 B.) 2,850.70 ft^3 C.) 2,980.27 ft^3 D.) 2,999.50 ft^3

OpenStudy (arabpride):

---> http://prntscr.com/52v6d9 <---

ganeshie8 (ganeshie8):

find the volume of small cone and multiply by the sf^3

ganeshie8 (ganeshie8):

sf = scale factor

OpenStudy (arabpride):

I know that the volume of the small cone is approximately 615.75ft^3

OpenStudy (arabpride):

So then my next step would be 615.75 x 3/4 ? 3/4 because it's the scale factor~

ganeshie8 (ganeshie8):

volume scales by sf^3

ganeshie8 (ganeshie8):

scale factor here is 20/12 right ?

OpenStudy (arabpride):

Oh that's right - cubed... yes it is~

OpenStudy (arabpride):

but i reduced it...

ganeshie8 (ganeshie8):

\[\text{volume of large cone} = \mathcal{(s.f)^3 \times }(\text{volume of small cone})\]

ganeshie8 (ganeshie8):

20/12 = (4*5)/(4*3) = 5/3 it reduces to 5/3, so sf would be 5/3

OpenStudy (arabpride):

OH - i reduced it wrong - i see... thanks cx so then i would multiply the volume of the smaller cone by 5/3 - correct?

OpenStudy (arabpride):

But I keep getting the same answer - 1026.25 ft^3 - But that isn't an option >.<

OpenStudy (anonymous):

(12/20)^3 = (pi/3)(7)^2(12) / V V = option B

OpenStudy (anonymous):

remember, in a similar objects, the volume ratio equals the cubed of the ratio of their sides

OpenStudy (anonymous):

and by sides, i meant the scale factor

OpenStudy (arabpride):

That makes sense, i was just confused because i used square instead of cube~ Thanks :3